BabyGurl:

Question : What is the derivative of f(x) = 15x^2 + 2x^2 + 3xf(x)=15x ​2 ​​ +2x ​2 ​​ +3x a) 34x + 3 b) -34x - 3 c) 34x - 3 d) 17x + 3

justjm:

is the question asking you to find derivative of \( f(x)=15x^2+2x^2+3x\) ? Combine like terms \( f(x) = 17x^2+3x\) Use power and sum rule \(f'(x)= \frac{d}{dx}17x^2 + \frac{d}{dx}3x \) \(f'(x)=34x+3\) Option A